OpenStudy (anonymous):

Find the solution of the square root of the quantity of x plus 2 plus 4 equals 6 and determine if it is an extraneous solution. x = 2; extraneous x = 2; not extraneous x = 30; extraneous x = 30; not extraneous

OpenStudy (imstuck):

\[\sqrt{(x+2)}+4=6\]

OpenStudy (imstuck):

is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

First this is to move the 4 to the other side, leaving this:

OpenStudy (imstuck):

\[\sqrt{(x+2)}=2\]and then square both sides to get rid of the radial sign, like this:

OpenStudy (imstuck):

\[(\sqrt{(x+2)})^{2}=(2)^{2}\]\[x+2=4\]

OpenStudy (imstuck):

x = 2. Is it extraneous?

OpenStudy (imstuck):

does it work or no?

OpenStudy (anonymous):

it works.. so its not extraneous correct? @IMStuck

OpenStudy (imstuck):

That is correct. Your answer is the second choice down.

OpenStudy (anonymous):

Thank you @IMStuck